# CREMA: Generalizable and Efficient Video-Language Reasoning via Multimodal Modular Fusion

### v2

@@ A.2 Multimodal Encoders and Visual Expert Models
We employ plug-and-play frozen experts to extract diverse modalities features, including depth map, optical flow, and surface normals, from raw videos. For optical flow estimation, we utilize the SotA Unimatch [67] model (GMFlow-scale2-regrefine6-mixdata). For depth map estimation, we leverage the SotA ZoeDepth-NK [4] model. For surface normals estimation, we follow Prismer [44] to use NLL-AngMF [3] that pre-trained on ScanNet [12]. We decode video into frames to extract per-frame depth map/optical flow/surface normals. We set 3 fps, 1 fps, and 3 fps to decode SQA3D, MUSIC-AVQA, and NExT-QA videos respectively.
@@ A.3 MMQA Pre-training Details
As discussed in Section 3.3, we conduct extra lightweight pre-training for MMQA module to obtain a good initialization. We follow audio-pertaining settings in X-InstructBLIP [49] with AudioCaps [31], but excluded caption data as our work is more focusing video reasoning. In this case, we obtained a QA-related subset from AudioCaps for MMQA-QA pre-training. Similarly, we utilized the 3D data released from 3D-LLM [20] and also took the QA format part for MMQA-3D pretraining. We pre-trained with $1e^{-5}$ learning rate and 1 epoch for efficient initialization.
@@ A.4 Baseline Model Implementation
We conduct experiments with 4 × 48GB A6000 GPUs, we report baseline model training hyperparameters in Table 9. We follow hyperparameter settings that have been searched to yield the best performance in SeViLA [69] with the same backbone model. To prompt LLM, we design different prompts for open-ended QA tasks (SQA3D, MUSIC-AVQA) and multi-choice QA (NExT-QA) following previous works [69, 18]. For open-ended QA, we let LLM generate responses without extra constraints and then compare the generated answers with ground-truth answers for accuracy calculation. We list prompts design for each dataset in Table 11.
@@ A.5 CREMA Implementation Details
CREMA framework adopts BLIP-2 [38], an image-language model with 4.1B parameters and pre-trained on 129M images in total, including COCO [40], Visual Genome [33], CC12M [55], SBU [48], and 115M images from LAION400M [54]. See Appendix for details. we also report our CREMA framwork training hyperparameters in Table 10. The experiments are conducted on the same 4 × 48GB A6000 GPUs machine.
@@
Mirasol3B TTM [51] V, A 66.4 3B
Mirasol3B [51] V, A 69.8 3B
@@ B.1 Extra Results on VGGSound and PerceptionTest
As listed in Table 12 and Table 13, we report extra fine-tuning performance on VGGSound [8] and PerceptionTest [50]. On VGGSound, we formulate the original audio-video classification task as an open-ended QA task. We calculate accuracy by directly matching the generated answer and the ground truth. Note that our original outputs have 2.9% (V) and 7.9% (V, A) predicted answers on test data that are out-of-vocabulary. We report both numbers with and without those out-of-vocabulary data in Section B.1. It shows that combined with audio (A) brings notable and consistent improvement. Our CREMA shows comparable performance with fewer trainable parameters compared with other methods. On PerceptionTest, our CREMA shows consistent improvement when adding new modalities while the baseline method struggles with this.
@@ B.4 The Effect of Modality-Sequential Training
We validate the effectiveness of the modality-sequential training of our approach in Table 16. We report the fine-tuning performance on three datasets: MUSIC-AVQA, SQA3D, and NExT-QA. As shown, our sequential training mechanism effectively improves multimodal video reasoning capabilities without additional architectures or the computationally and memory-intensive gradient modification strategy. This advantage is distinguishable from MLA [75], a multimodal representation learning method that sequentially trains a shared classification head on different modalities. Since MLA uses a unified trainable module (i.e., the shared classifier head) to train multimodal data, it employs the orthogonal gradient projection technique on different sensory inputs to mitigate the forgetting of previously learned modality information. In contrast, our model leverages substantial capacity to capture crucial information from various modalities through lightweight, yet effective modality-specific multi-query adapter (MMQA) modules, avoiding modality forgetting. Furthermore, we propose a modality-adaptive dynamic early exit strategy, enabling our model to converge quickly and mitigate over-convergence issues in different modality data by utilizing the gradients of lightweight modality-specific modules.
@@ Appendix C Limitation & Broader Impacts
Limitations. Our CREMA is based on BLIP-2 structure which was originally pre-trained only on image-language data. This may lead to limited temporal modeling/knowledge in the pre-trained model. It might not handle well fine-grained temporal events (e.g., open the door vs. close the door). Moreover, a potential limitation could be incorporating multimodal inputs such as depth and flow, which requires computing these modalities and introduces additional overhead.

### v4

@@ A.2 Multimodal Encoders and Visual Expert Models
We employ plug-and-play frozen experts to extract diverse modalities features, including depth map, optical flow, and surface normals, from raw videos. For optical flow estimation, we utilize the SotA Unimatch (Xu et al., 2023) model (GMFlow-scale2-regrefine6-mixdata). For depth map estimation, we leverage the SotA ZoeDepth-NK (Bhat et al., 2023) model. For surface normals estimation, we follow Prismer (Liu et al., 2023c) to use NLL-AngMF (Bae et al., 2021) that pre-trained on ScanNet (Dai et al., 2017). We decode video into frames to extract per-frame depth map/optical flow/surface normals. We set 3 fps, 1 fps, and 3 fps to decode SQA3D, MUSIC-AVQA, and NExT-QA videos respectively.
@@ A.3 MMQA Pre-training Details
As discussed in Section 3.3, we conduct extra lightweight pre-training for MMQA module to obtain a good initialization. We follow audio-pertaining settings in X-InstructBLIP (Panagopoulou et al., 2023) with AudioCaps (Kim et al., 2019), but excluded caption data as our work is more focusing video reasoning. In this case, we obtained a QA-related subset from AudioCaps for MMQA-QA pre-training. Similarly, we utilized the 3D data released from 3D-LLM (Hong et al., 2023) and also took the QA format part for MMQA-3D pretraining. We pre-trained with $1e^{-5}$ learning rate and 1 epoch for efficient initialization.
@@ A.4 Baseline Model Implementation
We conduct experiments with 4 × 48GB A6000 GPUs, we report baseline model training hyperparameters in Table 10. We follow hyperparameter settings that have been searched to yield the best performance in SeViLA (Yu et al., 2023a) with the same backbone model. To prompt LLM, we design different prompts for open-ended QA tasks (SQA3D, MUSIC-AVQA) and multi-choice QA (NExT-QA) following previous works (Yu et al., 2023a, Han et al., 2023). For open-ended QA, we let LLM generate responses without extra constraints and then compare the generated answers with ground-truth answers for accuracy calculation. We list prompts design for each dataset in Table 12. We utilize the same multimodal encoders, and multimodal information from the same estimator for fair comparison.
@@ A.5 CREMA Implementation Details
CREMA framework adopts BLIP-2 (Li et al., 2023b), an image-language model with 4.1B parameters and pre-trained on 129M images in total, including COCO (Lin et al., 2014), Visual Genome (Krishna et al., 2017), CC12M (Sharma et al., 2018), SBU (Ordonez et al., 2011), and 115M images from LAION400M (Schuhmann et al., 2021). See Appendix for details. we also report our CREMA framwork training hyperparameters in Table 11. The experiments are conducted on the same 4 × 48GB A6000 GPUs machine.
In the zero-shot setting, we conducted evaluations on SQA3D (video + point cloud) and MUSIC-AVQA (video + audio). Since these tests include only two modalities at a time, we bypassed the Self-Gated Multimodal Query Fusion module and directly concatenated the video tokens with the corresponding modality tokens. This ensures no parameter mismatch or interference and every modality is independent during inference.
@@
Mirasol3B TTM (Piergiovanni et al., 2023) V, A 66.4 3B
Mirasol3B (Piergiovanni et al., 2023) V, A 69.8 3B
@@ B.1 Extra Results on VGGSound and PerceptionTest
As listed in Table 13 and Table 14, we report extra fine-tuning performance on VGGSound (Chen et al., 2020) and PerceptionTest (Pătrăucean et al., 2023). On VGGSound, we formulate the original audio-video classification task as an open-ended QA task. We calculate accuracy by directly matching the generated answer and the ground truth. Note that our original outputs have 2.9% (V) and 7.9% (V, A) predicted answers on test data that are out-of-vocabulary. We report both numbers with and without those out-of-vocabulary data in Section B.1. It shows that combined with audio (A) brings notable and consistent improvement. Our CREMA shows comparable performance with fewer trainable parameters compared with other methods. On PerceptionTest, our CREMA shows consistent improvement when adding new modalities while the baseline method struggles with this.
@@ B.4 The Effect of Modality-Sequential Training
We validate the effectiveness of the modality-sequential training of our approach in Table 17. We report the fine-tuning performance on three datasets: MUSIC-AVQA, SQA3D, and NExT-QA. As shown, our sequential training mechanism effectively improves multimodal video reasoning capabilities without additional architectures or the computationally and memory-intensive gradient modification strategy. This advantage is distinguishable from MLA (Zhang et al., 2024), a multimodal representation learning method that sequentially trains a shared classification head on different modalities. Since MLA uses a unified trainable module (i.e., the shared classifier head) to train multimodal data, it employs the orthogonal gradient projection technique on different sensory inputs to mitigate the forgetting of previously learned modality information. In contrast, our model leverages substantial capacity to capture crucial information from various modalities through lightweight, yet effective modality-specific multi-query adapter (MMQA) modules, avoiding modality forgetting. Furthermore, we propose a modality-adaptive dynamic early exit strategy, enabling our model to converge quickly and mitigate over-convergence issues in different modality data by utilizing the gradients of lightweight modality-specific modules.
@@ B.7 Qualitative Analysis
Figure 5: Visualization on attention map under different modality combinations. Top: with audio and video. Bottom: with audio, optical flow, and video. We omit audio for simplicity. We highlight attention regions that may affect model prediction with red boxes.
B.8 Extra Discussions on Model Design.
Prioritizing different modalities. We conducted additional experiments comparing different prioritization strategies, including prioritizing other modalities and treating all modalities equally (i.e., no fusion, directly concatenating tokens) in Table 20. It shows that prioritizing video achieves the best performance, and prioritizing other modalities or treating all modalities equally results in lower performance, validating the effectiveness of our design.
Single FC layer Design. we conducted additional experiments comparing the FC layer to a lighter-weight alternative, interpolation to project dimension-mismatched features between multimodal encoder and multimodal Q-Former in Table 21. The results show that the FC layer improves performance while adding a negligible number of parameters. We attribute this improvement to the FC layer’s ability to provide a more dynamic and learnable projection between multimodal encoders and the Multimodal QFormer, which better aligns features compared to interpolation.
Query Token Length. We conducted experiments analyzing the impact of query token length on performance, trainable parameters, and computational cost in Table 22. We find increasing the number of query tokens indeed improves accuracy as more fine-grained features are captured. However, it also leads to increasing computational costs (GFLOPs). We find that 32 query tokens per frame strike a good balance between performance and efficiency. This design aligns with BLIP-2, ensuring strong performance without excessive computational overhead.
Setting NExT-QA
Major: V + Supportive: D,F,N 73.9
Major: D + Supportive: V,F,N 62.1
No Prioritizing: V, D, F, N 73.5
Table 20: Ablation on prioritizing different modalities.
Connector # Parameters MUSIC-AVQA
One-layer FC 0.3M 79.4
Interpolation 0M 78.9
Table 21: Ablation on the projection layer between multimodal encoder and multimodal Q-Former.
Modalities # Quey Token NExT-QA Trainable Param. GFlops
V 16 70.8 $\sim$4M 1.0K
V 32 71.6 $\sim$4M 1.3 K
V 64 72.0 $\sim$4M 2.1 K
V,F 16 71.8 $\sim$8M 1.4K
V,F 32 72.4 $\sim$8M 2.2K
V,F 64 72.9 $\sim$8M 6.2K
Table 22: Ablation on the number of query tokens.
@@ Appendix C Limitation & Broader Impacts
Limitations. Our CREMA is mainly based on BLIP-2 structure which was originally pre-trained only on image-language data. This may lead to limited temporal modeling/knowledge in the pre-trained model. It might not handle well fine-grained temporal events (e.g., open the door vs. close the door). Moreover, a potential limitation could be incorporating multimodal inputs such as depth and flow, which requires computing these modalities and introduces additional overhead.